# Creation of ultracold heteronuclear -wave Feshbach molecules

### v1

@@ Acknowledgements.
- [1] Note: See supplementary material. Cited by: Creation of ultracold heteronuclear -wave Feshbach molecules, Creation of ultracold heteronuclear -wave Feshbach molecules, Creation of ultracold heteronuclear -wave Feshbach molecules, Creation of ultracold heteronuclear -wave Feshbach molecules.
@@
- [12] (2020-08) Feshbach resonances in -wave three-body recombination within fermi-fermi mixtures of open-shell and closed-shell atoms. Phys. Rev. X 10, pp. 031037. External Links: Document, Link Cited by: §II, §II.
@@
- [16] (2021-09) Lee-Huang-Yang effects in the ultracold mixture of and with attractive interspecies interactions. Phys. Rev. Res. 3, pp. 033247. External Links: Document, Link Cited by: §IV.
- [17] (2022-02) Improved characterization of Feshbach resonances and interaction potentials between and atoms. Phys. Rev. A 105, pp. 023313. External Links: Document, Link Cited by: §III, Creation of ultracold heteronuclear -wave Feshbach molecules, Creation of ultracold heteronuclear -wave Feshbach molecules, Creation of ultracold heteronuclear -wave Feshbach molecules.
@@
- [21] (2019) Molscat: a program for non-reactive quantum scattering calculations on atomic and molecular collisions. Comput. Phys. Commun. 241, pp. 9–18. External Links: ISSN 0010-4655, Document, Link Cited by: §III, Creation of ultracold heteronuclear -wave Feshbach molecules, Creation of ultracold heteronuclear -wave Feshbach molecules.
@@
- [29] (2015-11) Emergence of chaotic scattering in ultracold er and dy. Phys. Rev. X 5, pp. 041029. External Links: Document, Link Cited by: §II, §II.
@@
I Supplemental Material
II Section S1: TEMPERATURE-DEPENDENT ATOM-LOSS SPECTRA
@@
Under the condition , the Lorentzian function in Eq. (IIS2) can be approximated as a delta function peaked at . Substituting this into Eq. (IIS1) and averaging the three-particle recombination rate with a Maxwell-Boltzmann distribution over the relative three-body collision energy , the position of maximum atom loss correlated with temperature follows:
@@
III Section S2: Calculation of
Theoretical molecular binding energies were calculated using the molscat package [21], with model parameters adapted from a newly calibrated molecular potential [17]. By varying the relative kinetic energy between atoms in the open channel, we simulated the scattering cross-section as a function of the magnetic field to determine resonance locations. The calculated results for and resonances are summarized in Fig. IIIS1. Linear functions are used to fit the correlation between resonance shifts and relative kinetic energies, allowing for the extraction of the theoretical magnetic moment difference between the open and closed channels.
IV Section S3: Fast magnetic field control for creating molecules
The primary magnetic field is provided by a pair of large coils. Due to limited feedback control bandwidth and Eddy current, the field typically needs over 10 ms to stabilize within 10 mG following a ramp. This residual fluctuation affects both the MA, especially in creating the pure FMs because of the small separation between the peaks, and the effective state transfer of the residual atoms to the state for their rapid removal, which is necessary to increase the lifetime of the FMs. To mitigate this issue, we implemented a pair of fast-switching coils [16] and use it to actively compensate the magnetic field fluctuations using feed-forward control. An example of the magnetic field feed-forward control sequence is shown in Fig. IVS2. This technique has allowed for the complete removal of residual atoms using microwave and blast pulses within 300 s after the main coil sweep.

### v2

@@ Acknowledgements.
- [1] Note: See supplementary material, which includes Refs. [32,34]. Cited by: Creation of ultracold heteronuclear -wave Feshbach molecules, Creation of ultracold heteronuclear -wave Feshbach molecules, Creation of ultracold heteronuclear -wave Feshbach molecules, Creation of ultracold heteronuclear -wave Feshbach molecules.
@@
- [12] (2020-08) Feshbach resonances in -wave three-body recombination within fermi-fermi mixtures of open-shell and closed-shell atoms. Phys. Rev. X 10, pp. 031037. External Links: Document, Link Cited by: §I, §I.
@@
- [16] (2021-09) Lee-huang-yang effects in the ultracold mixture of and with attractive interspecies interactions. Phys. Rev. Res. 3, pp. 033247. External Links: Document, Link Cited by: §III.
- [17] (2022-02) Improved characterization of Feshbach resonances and interaction potentials between and atoms. Phys. Rev. A 105, pp. 023313. External Links: Document, Link Cited by: §II, Creation of ultracold heteronuclear -wave Feshbach molecules, Creation of ultracold heteronuclear -wave Feshbach molecules, Creation of ultracold heteronuclear -wave Feshbach molecules.
@@
- [21] (2019) Molscat: a program for non-reactive quantum scattering calculations on atomic and molecular collisions. Comput. Phys. Commun. 241, pp. 9–18. External Links: ISSN 0010-4655, Document, Link Cited by: §II, Creation of ultracold heteronuclear -wave Feshbach molecules, Creation of ultracold heteronuclear -wave Feshbach molecules.
@@
- [29] (2015-11) Emergence of chaotic scattering in ultracold er and dy. Phys. Rev. X 5, pp. 041029. External Links: Document, Link Cited by: §I, §I.
@@
Supplemental Material
I Section S1: TEMPERATURE-DEPENDENT ATOM-LOSS SPECTRA
@@
Under the condition , the Lorentzian function in Eq. (IS2) can be approximated as a delta function peaked at . Substituting this into Eq. (IS1) and averaging the three-particle recombination rate with a Maxwell-Boltzmann distribution over the relative three-body collision energy , the position of maximum atom loss correlated with temperature follows:
@@
II Section S2: Calculation of
Theoretical molecular binding energies were calculated using the molscat package [21], with model parameters adapted from a newly calibrated molecular potential [17]. By varying the relative kinetic energy between atoms in the open channel, we simulated the scattering cross-section as a function of the magnetic field to determine resonance locations. The calculated results for and resonances are summarized in Fig. IIS1. Linear functions are used to fit the correlation between resonance shifts and relative kinetic energies, allowing for the extraction of the theoretical magnetic moment difference between the open and closed channels.
III Section S3: Fast magnetic field control for creating molecules
The primary magnetic field is provided by a pair of large coils. Due to limited feedback control bandwidth and Eddy current, the field typically needs over 10 ms to stabilize within 10 mG following a ramp. This residual fluctuation affects both the MA, especially in creating the pure FMs because of the small separation between the peaks, and the effective state transfer of the residual atoms to the state for their rapid removal, which is necessary to increase the lifetime of the FMs. To mitigate this issue, we implemented a pair of fast-switching coils [16] and use it to actively compensate the magnetic field fluctuations using feed-forward control. An example of the magnetic field feed-forward control sequence is shown in Fig. IIIS2. This technique has allowed for the complete removal of residual atoms using microwave and blast pulses within 300 s after the main coil sweep.